# Think before you speak: Training Language Models With Pause Tokens

### v2

@@ Abstract
Transformer-based language models generate responses by producing a series of tokens in immediate succession: the $(K+1)^{\rm th}$ token is an outcome of manipulating $K$ hidden vectors per layer, one vector per preceding token. What if instead we were to let the model manipulate say, $K+10$ hidden vectors, before it outputs the $(K+1)^{\rm th}$ token? We operationalize this idea by performing training and inference on language models with a (learnable) pause token, a sequence of which is appended to the input prefix. We then delay extracting the model’s outputs until the last pause token is seen, thereby allowing the model to process extra computation before committing to an answer. We empirically evaluate pause-training on decoder-only models of 1B and 130M parameters with causal pretraining on C4, and on downstream tasks covering reasoning, question-answering, general understanding and fact recall. Our main finding is that inference-time delays show gains on our tasks when the model is both pre-trained and finetuned with delays. For the 1B model, we witness gains on eight tasks, most prominently, a gain of $18\%$ EM score on the QA task of SQuAD, $8\%$ on CommonSenseQA and $1\%$ accuracy on the reasoning task of GSM8k. Our work raises a range of conceptual and practical future research questions on making delayed next-token prediction a widely applicable new paradigm.
@@ 1 Introduction
    We pose the question of what happens if we delay a model’s answer generation, and how can we execute these delays? We design one way: training with dummy <pause> tokens. Accordingly, we design a pause-injected pretraining, downstream finetuning and inference procedure.
@@ 2 Preliminaries
We briefly outline the next-token prediction process in a standard causal decoder-only language model (details in §A). Consider a vocabulary $\mathcal{V}$ and an input $\mathbf{p}_{1:K}\in\mathcal{V}^{K}$ of $K$ tokens. Let $f$ to denote a Transformer-based language model, from which we sample the next token as $p_{K+1}\sim f(\mathbf{p}_{1:K})$. To achieve this, internally, each layer $l\in[1,L]$ of the Transformer produces an intermediate vector $\mathbf{v}_{k}^{(l)}$ corresponding to each input token. The next token i.e. $p_{K+1}$ is then sampled from a distribution inferred from the last vector in the last layer, $\mathbf{v}^{(L)}_{K}$.
On a high level, each layer in the above process can be represented as a function $T:\mathbb{R}^{D\times K}\to\mathbb{R}^{D\times K}$. Its input is a matrix of $K$ vectors, $V_{1:K}=\left[\mathbf{v}_{1},\ldots,\mathbf{v}_{K}\right]$, and likewise, the output, $V^{\prime}_{1:K}$. This mapping itself involves two key (parameterized) modules. The first is the attention module $\Phi_{\rm Attn}$ which takes as inputs two matrices $V_{\rm query},V_{\rm value}\in\mathbb{R}^{D\times N}$ (for any $N$) and a “value” vector $\mathbf{v}\in\mathbb{R}^{D}$ to produce an output vector in $\mathbb{R}^{D}$. This is followed by a feedforward module $f_{\rm FF}:\mathbb{R}^{D}\to\mathbb{R}^{D}$. Then, given the inputs $\mathbf{v}_{k}$, and given the layer-norm module $\Phi_{\rm LN}:\mathbb{R}^{D}\to\mathbb{R}^{D}$, the outputs $\mathbf{v}_{k}^{\prime}$ for $k\leq K$ can be expressed as:
@@ 3.1 Learning and inference with the <pause> token
A simple choice for the dummy tokens are special characters such as ‘.‘ or ‘#‘, as Lanham et al. (2023) chose for inference. But to prevent the model from confounding the role of delays with the role the above characters play in natural language, we choose a single <pause> token residing outside of the standard vocabulary. To impose multi-token delays, we simply repeat this token. Building on this core idea, below we discuss our specific techniques for pause-pretraining and pause-finetuning.
@@ 8 Conclusion, Limitations and Future Work
Acknowledgements: We would like to thank Kaifeng Lyu, Srinadh Bhojanapalli, Yongchao Zhou, Nikunj Saunshi and Seungyeon Kim for helping set up the initial codebase and for their guidance.
@@ Appendix C Prepending vs Appending Pause Tokens
Dataset Metric StdPTStdFT StdPTPauseFT PausePTPauseFT
@@ Appendix H Hyperparameters: Downstream finetuning
    $\textbf{p}_{1:N}\sim\mathcal{D}_{\rm pt}$  /* Input Sequence from corpus */
    $\tilde{\textbf{p}}_{1:N+M_{pt}}=\texttt{random\_insert}(\textbf{p}_{1:N},\texttt{<pause>}{},M_{\rm pt})$  /* Insert $M_{\rm pt}$ pause tokens randomly in the original input sequence $\textbf{p}_{1:N}$, extending its length by $M_{\rm pt}$/
@@
    $\mathcal{L}_{\rm PausePT}(f_{\theta},\tilde{\mathbf{p}}_{1:N+M_{\rm pt}})=\sum_{k=1,k\notin S_{\rm ignore}}^{N+M_{\rm pt}-1}\mathcal{L}_{\rm CE}(\tilde{p}_{k+1},f_{\theta}(\tilde{\textbf{p}}_{1:k}))$   /* Next token prediction error excludes targets which are pause (model isn’t made to learn to predict pause itself) */
    $\theta=\theta-\nabla_{\theta}\mathcal{L}_{\rm PausePT}(f_{\theta},\tilde{\mathbf{p}}_{1:N+M_{\rm pt}})$   /* Update the model */
@@
    $\textbf{p}_{1:N},\textbf{t}_{1:T}\sim\mathcal{D}_{\rm ft}$  /* Sample prefix and target */
    $\tilde{\textbf{p}}_{1:N+M_{\text{ft}}}=\text{Concat}[\textbf{p}_{1:N},[\texttt{<pause>}{}]\times M_{\text{ft}}]$  /* Append prefix and $M_{\rm ft}$ pauses */
    $\mathcal{L}_{\rm PauseFT}(f_{\theta},\tilde{\textbf{p}}_{1:N+M_{\text{ft}}},\textbf{t}_{1:T})=\sum_{k=0}^{T-1}\mathcal{L}_{\text{CE}}(t_{k+1},f_{\theta}(\text{Concat}[\tilde{\textbf{p}}_{1:N+M_{\rm ft}},\textbf{t}_{1:k}])$   /* Next token prediction error on targets */
@@
    $\tilde{\textbf{p}}_{[1:N+M_{\text{inf}}]}=[\textbf{p}_{1:N},[\texttt{<pause>}{}]\times M_{\text{inf}}]$  /* Append $M_{\rm inf}$ pauses to prefix */
    $\tilde{\textbf{p}}_{N+M_{\text{inf}}+1}\sim f_{\theta}(\tilde{\textbf{p}}_{1:N+M_{\text{inf}}})$   /* Predict the next token in the sequence, and continue in an auto-regressive fashion */
@@ I.2.1 Supporting lemmas for estimating FLOPS efficiency
(structure of the underlying task) Given the vocabulary space $\mathcal{V}$, let $\circ$ be a generic 2-ary operator on $\mathcal{V}$ For an input sequence length $L$, consider a corresponding function class $\mathcal{F}_{L}$ that corresponds to all functions $f:\mathcal{V}^{L}\mapsto\mathcal{V}$ that require applying $N$circ operations independently following by a generic aggregation operation $g_{\rm aggr}:\mathcal{V}^{N}\to\mathcal{V}$:
$\mathcal{F}_{L}$ =$=\Bigg{\{}f:\mathcal{V}^{L}\mapsto\mathcal{V}\Bigg{|}\,\exists i_{1},\ldots i_{N},j_{1},\ldots j_{N}\in[1,L],$ (18)
[…]

### v3

@@ Abstract
Transformer-based language models generate responses by producing a series of tokens in immediate succession: the $(K+1)^{\rm th}$ token is an outcome of manipulating $K$ hidden vectors per layer, one vector per preceding token. What if instead we were to let the model manipulate say, $K+10$ hidden vectors, before it outputs the $(K+1)^{\rm th}$ token? We operationalize this idea by performing training and inference on language models with a (learnable) pause token, a sequence of which is appended to the input prefix. We then delay extracting the model’s outputs until the last pause token is seen, thereby allowing the model to process extra computation before committing to an answer. We empirically evaluate pause-training on decoder-only models of 1B and 130M parameters with causal pretraining on C4, and on downstream tasks covering reasoning, question-answering, general understanding and fact recall. Our main finding is that inference-time delays show gains on our tasks when the model is both pretrained and finetuned with delays. For the 1B model, we witness gains on eight tasks, most prominently, a gain of $18\%$ EM score on the QA task of SQuAD, $8\%$ on CommonSenseQA and $1\%$ accuracy on the reasoning task of GSM8k. Our work raises a range of conceptual and practical future research questions on making delayed next-token prediction a widely applicable new paradigm.
@@ 1 Introduction
    We pose the question of what happens if we delay a model’s answer generation, and how can we execute these delays? We design one way: training with dummy <pause> tokens. Accordingly, we design a pause-injected pretraining, downstream finetuning, and inference procedure.
@@ 2 Preliminaries
We briefly outline the next-token prediction process in a standard causal decoder-only language model (details in §A). Consider a vocabulary $\mathcal{V}$ and an input $\mathbf{p}_{1:K}\in\mathcal{V}^{K}$ of $K$ tokens. Let $f$ denote a Transformer-based language model, from which we sample the next token as $p_{K+1}\sim f(\mathbf{p}_{1:K})$. To achieve this, internally, each layer $l\in[1,L]$ of the Transformer produces an intermediate vector $\mathbf{v}_{k}^{(l)}$ corresponding to each input token. The next token, i.e., $p_{K+1}$ is then sampled from a distribution inferred from the last vector in the last layer, $\mathbf{v}^{(L)}_{K}$.
On a high level, each layer in the above process can be represented as a function $T:\mathbb{R}^{D\times K}\to\mathbb{R}^{D\times K}$. Its input is a matrix of $K$ vectors, $V_{1:K}=\left[\mathbf{v}_{1},\ldots,\mathbf{v}_{K}\right]$, and likewise, the output, $V^{\prime}_{1:K}$. This mapping itself involves two key (parameterized) modules. The first is the attention module $\Phi_{\rm Attn}$ which takes as inputs two matrices $V_{\rm key},V_{\rm value}\in\mathbb{R}^{D\times N}$ (for any $N$) and a “query” vector $\mathbf{v}\in\mathbb{R}^{D}$ to produce an output vector in $\mathbb{R}^{D}$. This is followed by a feedforward module $f_{\rm FF}:\mathbb{R}^{D}\to\mathbb{R}^{D}$. Then, given the inputs $\mathbf{v}_{k}$, and given the layer-norm module $\Phi_{\rm LN}:\mathbb{R}^{D}\to\mathbb{R}^{D}$, the outputs $\mathbf{v}_{k}^{\prime}$ for $k\leq K$ can be expressed as:
@@ 3.1 Learning and inference with the <pause> token
A simple choice for the dummy tokens are special characters such as ‘.’ or ‘#’, as Lanham et al. (2023) chose for inference. But to prevent the model from confounding the role of delays with the role the above characters play in natural language, we choose a single <pause> token residing outside of the standard vocabulary. To impose multi-token delays, we simply repeat this token. Building on this core idea, below we discuss our specific techniques for pause-pretraining and pause-finetuning.
@@ 8 Conclusion, Limitations and Future Work
Acknowledgements: We would like to thank Kaifeng Lyu, Srinadh Bhojanapalli, Yongchao Zhou, Nikunj Saunshi, and Seungyeon Kim for helping set up the initial codebase and for their guidance.
@@ Appendix C Prepending vs Appending Pause Tokens
Dataset Metric StdPT_StdFT StdPT_PauseFT PausePT_PauseFT
@@ Appendix H Hyperparameters: Downstream finetuning
    $\textbf{p}_{1:N}\sim\mathcal{D}_{\rm pt}$
  /* Input Sequence from corpus */
    $\tilde{\textbf{p}}_{1:N+M_{pt}}=\texttt{random\_insert}(\textbf{p}_{1:N},\texttt{<pause>}{},M_{\rm pt})$
  /* Insert $M_{\rm pt}$ pause tokens randomly in the original input sequence $\textbf{p}_{1:N}$, extending its length by $M_{\rm pt}$/
@@
  /* Identify the set of positions where the next token is <pause> */
    $\mathcal{L}_{\rm PausePT}(f_{\theta},\tilde{\mathbf{p}}_{1:N+M_{\rm pt}})=\sum_{k=1,k\notin S_{\rm ignore}}^{N+M_{\rm pt}-1}\mathcal{L}_{\rm CE}(\tilde{p}_{k+1},f_{\theta}(\tilde{\textbf{p}}_{1:k}))$
  /* Next token prediction error excludes targets which are pause (model isn’t made to learn to predict pause itself) */
    $\theta=\theta-\nabla_{\theta}\mathcal{L}_{\rm PausePT}(f_{\theta},\tilde{\mathbf{p}}_{1:N+M_{\rm pt}})$
  /* Update the model */
@@
    $\textbf{p}_{1:N},\textbf{t}_{1:T}\sim\mathcal{D}_{\rm ft}$
  /* Sample prefix and target */
    $\tilde{\textbf{p}}_{1:N+M_{\text{ft}}}=\text{Concat}[\textbf{p}_{1:N},[\texttt{<pause>}{}]\times M_{\text{ft}}]$
  /* Append prefix and $M_{\rm ft}$ pauses */
    $\mathcal{L}_{\rm PauseFT}(f_{\theta},\tilde{\textbf{p}}_{1:N+M_{\text{ft}}},\textbf{t}_{1:T})=\sum_{k=0}^{T-1}\mathcal{L}_{\text{CE}}(t_{k+1},f_{\theta}(\text{Concat}[\tilde{\textbf{p}}_{1:N+M_{\rm ft}},\textbf{t}_{1:k}])$
  /* Next token prediction error on targets */
@@
    $\tilde{\textbf{p}}_{[1:N+M_{\text{inf}}]}=[\textbf{p}_{1:N},[\texttt{<pause>}{}]\times M_{\text{inf}}]$
  /* Append $M_{\rm inf}$ pauses to prefix */
    $\tilde{\textbf{p}}_{N+M_{\text{inf}}+1}\sim f_{\theta}(\tilde{\textbf{p}}_{1:N+M_{\text{inf}}})$
  /* Predict the next token in the sequence, and continue in an auto-regressive fashion */
@@ I.2.1 Supporting lemmas for estimating FLOPS efficiency
(structure of the underlying task) Given the vocabulary space $\mathcal{V}$, let $\circ$ be a generic 2-ary operator on $\mathcal{V}$ For an input sequence length $L$, consider a corresponding function class $\mathcal{F}_{N}$ that corresponds to all functions $f:\mathcal{V}^{L}\mapsto\mathcal{V}$ that require applying $N$circ operations independently following by a generic aggregation operation $g_{\rm aggr}:\mathcal{V}^{N}\to\mathcal{V}$:
$\mathcal{F}_{N}$ =$=\Bigg{\{}f:\mathcal{V}^{L}\mapsto\mathcal{V}\Bigg{|}\,\exists i_{1},\ldots i_{N},j_{1},\ldots j_{N}\in[1,L],$ (18)
$\left.f(v_{1},v_{2},\ldots,v_{L})=g_{\text{aggr}}\big{(}\underbrace{(v_{i_{1}}\circ v_{j_{1}}),(v_{i_{2}}\circ v_{j_{2}}),\ldots,(v_{i_{N}}\circ v_{j_{N}})}_{N\text{independent }\circ\text{ operations.}}\right.\big{)}\Bigg{\}}.$ (19)